# Emergent chirality and superfluidity of parity-doubled baryons in neutron stars

### v1

@@ Abstract
We propose novel superfluids induced by the parity-doubled baryons. The parity-doubled baryons, i.e., a nucleon $N(940)$ with spin-parity $J^{P}=1/2^{+}$ and an excited nucleon $N^{\ast}(1535)$ with $J^{P}=1/2^{-}$ in vacuum, become degenerate at sufficiently high density where the chiral symmetry is restored. In this study, we extend the conventional $\mathrm{U}(1)$ chiral symmetry to the higher dimensional symmetries, dubbed emergent chiral symmetries, including the naive and mirror assignments as their subgroups. Starting with the Lagrangian up to four-point interactions among the neutron $n$ and its chiral partner $n^{\ast}$, neutral components in $N$ and $N^{\ast}$, in pure neutron matter, we investigate the properties of the ground state with a pairing gap generated by the $n$ and $n^{\ast}$ in the mean-field approximation. We find vector-type condensates that induce the dynamical breaking of a new class of internal symmetries, emergent chiral symmetries, as well as the baryon number and the rotational symmetries of the real space, indicating the appearance of massless Nambu-Goldstone bosons consisting of six quarks: emergent pions, superfluid phonons, and magnons, respectively. We also study the fermionic excitation modes at low-energy scales, and show that they exhibit a spatial anisotropy of the propagation at the Dirac cone in momentum space. Some phenomenological implications are advocated, shedding new light on the properties of neutron stars.
@@ I Introduction
Recent advances in multi-messenger astronomy have provided a key set of constraints on the equation of state (EoS) of QCD at high baryon density, via extraterrestrial observations of dense stellar objects such as neutron stars (NSs) [1]. It was suggested [2] that the softening of the EoS at intermediate densities required to comply with the observed tidal deformability, together with the subsequent stiffening at high densities to protect the $2M_{\odot}$ NSs from their gravitational collapse, may indicate a phase transition at the core. This is often attributed to color deconfinement [3, 4, 5], although it is rather a heuristic guidance within simplistic approaches that do not account for the microscopic dynamics of underlying QCD.
@@
Whereas the existence of quark matter is an intriguing option, a pure hadronic scenario has been shown to fully reconcile the modern constraints from observation [6, 7, 8, 9, 10, 11], where the essential ingredients are the lowest nucleon $N(940)$ and its chiral partner $N^{*}(1535)$ interacting with the scalar and pseudoscalar mesons in a chiral-invariant way. The properties of $N$ and $N^{*}$ can be well captured within the parity-doublet model based on the mirror assignment of chirality [12, 13]: The mirror assignment is a more general representation for the nucleons with opposite parity, than the conventional one à la Gell-Mann and Levy, the naive assignment. It is a striking difference that the parity doublet model predicts the $N$ and $N^{*}$ being degenerate and massive when the chiral symmetry becomes restored, whereas the naive one inevitably leads to massless nucleons. The former, the mirror scenario, emerges in Nature, as confirmed by the recent lattice QCD simulations exhibiting a clear manifestation of parity doubling near the chiral crossover in the baryon spectra [14, 15, 16]. Thus, the hadronic matter at high density has recently attracted attention as an interesting realization of QCD substances in the confined phase.
@@
At high density, nucleons may form Cooper pairs giving rise to superfluidity. Neutron superfluids have been studied intensively not only in astrophysics and nuclear physics but also in condensed matter physics [17, 18, 19, 20, 21]. It has been discussed since the early studies that the neutron ${}^{1}S_{0}$ (spin-singlet and $S$-wave) superfluids exist at low densities near the surface of neutron stars [22, 23], and the neutron ${}^{3}P_{2}$ (spin-triplet and $P$-wave) superfluids exist at higher densities in the deep inside [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. The neutron ${}^{3}P_{2}$ superfluids are tolerant against the strong magnetic fields such as in magnetars with $B=10^{15}$-$10^{18}$ G, thus are important to understand the interior of neutron stars and magnetars. In fact, it was studied that the enhancement of neutrino emission due to the neutron ${}^{3}P_{2}$ superfluids may explain the rapid cooling observed for the neutron star in Casiopeia A [55, 56, 57] (see also Refs. [58, 59, 60]). It was also discussed that the bosonic excitations in the neutron ${}^{3}P_{2}$ superfluids may be relevant for the cooling process by neutrino emission [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73]. It has been proposed that vortices penetrating ${}^{3}P_{2}$-${}^{1}S_{0}$ interfaces inside the neutron stars can explain pulsar glitches without fine tuning [53]. In the condensed-matter aspects, it is particularly interesting that the ${}^{3}P_{2}$ superfluids are topological superfluids, admitting topologically protected Majorana fermions on their surfaces [35] and inside cores [40, 47] of singly quantized vortices [32, 33, 34] and half-quantum vortices [37, 49, 50] (see Ref. [51] as a review), and thus they may be called the largest topological matter in universe.
@@
In the past studies, it was conventionally assumed that the chiral symmetry is spontaneously broken in the neutron ${}^{1}S_{0}$ and ${}^{3}P_{2}$ superfluids as in vacuum. However, it is expected that the chiral symmetry can become restored at sufficiently high density inside neutron stars even if the substances remain composed of hadrons, such as in the parity-doublet model. It is of particular interest to explore the properties of neutron superfluids embracing potential modifications of hadrons due to (partially) restored chiral symmetry.
@@
In this article, we propose superfluidity of parity-doubled neutrons for the first time. This new type of superfluid is characterized by the Cooper pairings given not only by $nn$ but also $nn^{\ast}$ and $n^{\ast}n^{\ast}$, where $n^{\ast}$ (spin-parity $J^{P}=1/2^{-}$) is an excited state of a neutron $n$, i.e., the neutral state of $N^{\ast}$. It is natural to introduce an internal symmetry for the two fermions in a doublet, in particular an exchange symmetry between them, since there is no distinction between them a priori; In addition to the conventional chiral symmetry, the doublet can have an extra symmetry, either $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ (the relative phase of the doublet) or $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$, dubbed an emergent chiral symmetry. The emergent chiral symmetries include the naive and mirror assignments as subgroups, thus providing us in-depth understanding of the chiral symmetry. We will construct the full Lagrangian including four-point interactions among the neutron $n$ and its chiral partner $n^{\ast}$ and present the phase structure in the mean field approximation. We find novel vector-type $J^{P}=1^{+}$ and $1^{-}$ condensates (${}^{3}P_{1}$ and ${}^{3}S_{1}$-${}^{3}D_{2}$ in nonrelativistic notation) exhibiting the dynamical breaking of the emergent chiral symmetries, the baryon number symmetry, and the rotational symmetry of the real space, resulting in emergent pions, superfluid phonons, and magnons, respectively as gapless Nambu-Goldstone (NG) bosons consisting of six quarks. We also find low-energy fermionic excitation modes exhibiting a spatial anisotropy of the propagation. We shall name those new states of superfluidity emergent chiral superfluids. The emergent chiral superfluids may exist in a more inner part of NSs than the phases of neutron ${}^{1}S_{0}$ and ${}^{3}P_{2}$ superfluids as illustrated in Fig. 1.
@@
The paper is organized as follows: In Sec. II, we introduce the emergent chiral symmetry for the parity-doubled neutrons as an extension of the usual chiral symmetry. Based on the emergent chiral symmetry, we construct the Lagrangain with four-point interactions for the chiral-doublet baryons. In Sec. III, we adopt the Namb-Gor’kov formalism to the particle-particle (hole-hole) interaction in the parity-doubled baryons and apply the mean field approximation by introducing gap functions. We show that the vector-type condensate accompanies the breaking of the emergent chiral symmetry, baryon number symmetry and rotational symmetry in the ground state. We also analyze the fermionic excitations in low-energy scale and show the existence of the Dirac points in the momentum space. The final section is devoted to our conclusion and outlook. In Appendix A, the naive and mirror assignments are briefly reviewed along with the emergent chiral symmetry. In Appendix B, the Lagrangians in the two assignments are given for the simplest emergent chiral symmetry. In Appendix C, the Fierz identities are summarized for convenience of readers. In Appendix D, the eigenfunctions of the Hamiltonian for the massless Dirac fermions are summarized.
@@
Figure 1: A schematic cross section inside a neutrons star formed by the crust, the neutron ${}^{1}S_{0}$ superfluids near the surface, the neutron ${}^{3}P_{2}$ superfluids at intermediate region, and the neutron emergent chiral superfluids in the deep interior. Phases of the neutron emergent chiral superfluids are denoted by ${\cal S}_{a}1^{-}$ (${}^{3}P_{1}$) ($a=0,1,3$) and ${\cal A}1^{+}$ (${}^{3}S_{1}$-${}^{3}D_{1}$). Here, ${\cal S}_{a}$ and ${\cal A}$ indicate the symmetric and antisymmetric representations in the emergent chiral symmetry, and $J^{P}$ represent total angular momentum $J$ and parity $P$. In the parentheses, the nonrelativistic quantum numbers, corresponding to $J^{P}=1^{-}$ and $1^{+}$, are denoted by ${}^{2S+1}L_{J}$ with spin $S$ and angular momentum $L$.
@@ II.1 Naive and mirror assignments
To describe consistently a system for baryons both in the broken and restored phases of chiral symmetry, we introduce parity-doubled baryons which are composed of a pair of a nucleon $N$ with spin-parity $J^{P}=1/2^{+}$ and an excited nucleon $N^{\ast}(1535)$ with $J^{P}=1/2^{-}$ [74]. The parity-doublet model has been applied to the hadron spectroscopy related to the chiral dynamics, see, e.g., Ref. [13] for a review. In this framework, the nucleon $N$ and the excited nucleon $N^{\ast}(1535)$ are expressed as a linear combination of two baryons denoted by $\Psi_{1}$ and $\Psi_{2}$ in the flavor eigenstates. Assigning the chiralities to $\Psi_{1}$ and $\Psi_{2}$ ($N$ and $N^{\ast}$), two different types of chiral symmetry have been considered: a naive assignment and a mirror assignment [74]. In the naive assignment, the left (right) component of $\Psi_{1}$ and the left (right) component of $\Psi_{2}$ rotates simultaneously. In the mirror assignment, in contrast, the left (right) component of $\Psi_{1}$ and the right (left) component of $\Psi_{2}$ rotates simultaneously. Notice the different combinations of the left and right components for $\Psi_{1}$ and $\Psi_{2}$, respectively, in the mirror assignment. 111 A parity doublet has also been referred to as a chiral doublet in the literature, e.g. [75, 76, 77, 78].
To model a neutron superfluid in the context of parity doublet framework, we shall consider the interaction between two parity-doubled baryons, i.e., $N$ and $N^{\ast}$. Because our main interest is primarily put on the neutron stars, we consider only the neutral components $n$ and $n^{\ast}$ by neglecting the charged ones. Here $n^{\ast}$ indicates $n^{\ast}(1535)$, i.e., the neutral component of $N^{\ast}(1535)$, for shorthand notation. Thus, $n$ and $n^{\ast}$ are expressed as a linear combination of $\psi_{1}$ and $\psi_{2}$: $n=\cos\theta\psi_{1}+\gamma_{5}\sin\theta\psi_{2}$ and $n^{\ast}=-\gamma_{5}\sin\theta\psi_{1}+\cos\theta\psi_{2}$ with an appropriate mixing angle $\theta$ [12, 13]. Here we consider that $\psi_{1}$ and $\psi_{2}$ have only neutral components in $\Psi_{1}$ and $\Psi_{2}$, respectively. Characterising our superfluid, we have $nn$, $nn^{\ast}$ and $n^{\ast}n^{\ast}$ interactions to get the particle-particle (hole-hole) pairings. We will consider a full set of four-point interactions (zero-range force) for $n$ and $n^{\ast}$. This setup is much different from the previous treatments of the parity-doublet model [6, 7, 8, 9, 10], where the $\sigma$ and $\omega$ fields are introduced to couple to the parity-doubled baryons like in the relativistic mean-field model of nuclear matter [79].

### v2

@@ Abstract
We propose novel superfluids induced by the parity-doubled baryons. The parity-doubled baryons, i.e., a nucleon $N(940)$ with mass 940 MeV and spin-parity $J^{P}=1/2^{+}$ and an excited nucleon $N^{\ast}(1535)$ with mass 1535 MeV and spin-parity $J^{P}=1/2^{-}$ in vacuum, become degenerate at sufficiently high density where the chiral symmetry is restored. In this study, we extend the conventional $\mathrm{U}(1)$ chiral symmetry to the higher dimensional symmetries, dubbed emergent chiral symmetries, including the naive and mirror assignments as their subgroups. Starting with the Lagrangian up to four-point interactions among the neutron $n$ and its chiral partner $n^{\ast}$, neutral components in $N$ and $N^{\ast}$, in pure neutron matter, we investigate the properties of the ground state with a pairing gap generated by the $n$ and $n^{\ast}$ in the mean-field approximation. We find vector-type condensates that induce the dynamical breaking of a new class of internal symmetries, emergent chiral symmetries, as well as the baryon number and the rotational symmetries of the real space, indicating the appearance of massless Nambu-Goldstone bosons consisting of six quarks: emergent pions, superfluid phonons, and magnons, respectively. We also study the fermionic excitation modes at low-energy scales, and show that they exhibit a spatial anisotropy of the propagation at the Dirac cone in momentum space. Some phenomenological implications are advocated, shedding new light on the properties of neutron stars.
@@ I Introduction
Recent advances in multimessenger astronomy have provided a key set of constraints on the equation of state (EoS) of QCD at high baryon density, via extraterrestrial observations of dense stellar objects such as neutron stars (NSs) [1]. It was suggested [2] that the softening of the EoS at intermediate densities required to comply with the observed tidal deformability, together with the subsequent stiffening at high densities to protect the $2M_{\odot}$ NSs from their gravitational collapse, may indicate a phase transition at the core. This is often attributed to color deconfinement [3, 4, 5], although it is rather a heuristic guidance within simplistic approaches that do not account for the microscopic dynamics of underlying QCD.
@@
Whereas the existence of quark matter is an intriguing option, a pure hadronic scenario has been shown to fully reconcile the modern constraints from observation [6, 7, 8, 9, 10, 11], where the essential ingredients are the lowest nucleon $N(940)$ (mass 940 MeV and spin-parity $J^{P}=1/2^{+}$) and its chiral partner $N^{*}(1535)$ (mass 1535 MeV and spin-parity $J^{P}=1/2^{-}$) interacting with the scalar and pseudoscalar mesons in a chiral-invariant way. The properties of $N$ and $N^{*}$ can be well captured within the parity-doublet model based on the mirror assignment of chirality [12, 13]: the mirror assignment is a more general representation for the nucleons with opposite parity, than the conventional one à la Gell-Mann and Levy, the naive assignment. It is a striking difference that the parity doublet model predicts the $N$ and $N^{*}$ being degenerate and massive when the chiral symmetry becomes restored, whereas the naive one inevitably leads to massless nucleons. The former, the mirror scenario, emerges in nature, as confirmed by the recent lattice QCD simulations exhibiting a clear manifestation of parity doubling near the chiral crossover in the baryon spectra [14, 15, 16]. Thus, the hadronic matter at high density has recently attracted attention as an interesting realization of QCD substances in the confined phase (see, e.g., a recent review Ref. [17]).
@@
At high density, nucleons may form Cooper pairs giving rise to superfluidity. Neutron superfluids have been studied intensively not only in astrophysics and nuclear physics but also in condensed matter physics [18, 19, 20, 21, 22]. It has been discussed since the early studies that the neutron ${}^{1}S_{0}$ (spin-singlet and $S$-wave) superfluids exist at low densities near the surface of neutron stars [23, 24], and the neutron ${}^{3}P_{2}$ (spin-triplet and $P$-wave) superfluids exist at higher densities in the deep inside [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. The neutron ${}^{3}P_{2}$ superfluids are tolerant against the strong magnetic fields such as in magnetars with $B=10^{15}$-$10^{18}$ G thus are important to understand the interior of neutron stars and magnetars. In fact, it was studied that the enhancement of neutrino emission due to the neutron ${}^{3}P_{2}$ superfluids may explain the rapid cooling observed for the neutron star in Casiopeia A [56, 57, 58] (see also Refs. [59, 60, 61]). It was also discussed that the bosonic excitations in the neutron ${}^{3}P_{2}$ superfluids may be relevant for the cooling process by neutrino emission [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74]. It has been proposed that vortices penetrating ${}^{3}P_{2}$-${}^{1}S_{0}$ interfaces inside the neutron stars can explain pulsar glitches without fine-tuning [54] (see also, e.g., a recent review Ref. [75]). [^1] It has been discussed that, beyond glitches, superfluidity might also influence pulsar timing noise in neutron stars [76]. In the condensed-matter aspects, it is particularly interesting that the ${}^{3}P_{2}$ superfluids are topological superfluids, admitting topologically protected Majorana fermions on their surfaces [36] and inside cores [41, 48] of singly quantized vortices [33, 34, 35] and half-quantum vortices [38, 50, 51] (see Ref. [52] as a review), and thus they may be called the largest topological matter in the Universe.
@@
In the past studies, it was conventionally assumed that the chiral symmetry is spontaneously broken in the neutron ${}^{1}S_{0}$ and ${}^{3}P_{2}$ superfluids as in vacuum (see, e.g., a review Ref. [20]). However, it is expected that the chiral symmetry can become restored at sufficiently high density inside neutron stars even if the substances remain composed of hadrons, such as in the parity-doublet model. For an example, it has been discussed that the baryon number density of hadronic matter could reach maximally $3$-$5n_{0}$ ($n_{0}=0.16$ fm-3: normal nuclear matter density) (see, e.g., Refs. [6, 7, 8] and references therein). It is of particular interest to explore the properties of neutron superfluids embracing potential modifications of hadrons due to (partially) restored chiral symmetry.
@@
In this article, we propose superfluidity of parity-doubled neutrons for the first time. This new type of superfluid is characterized by the Cooper pairings given not only by $nn$ but also $nn^{\ast}$ and $n^{\ast}n^{\ast}$, where $n^{\ast}$ (spin-parity $J^{P}=1/2^{-}$) is an excited state of a neutron $n$, i.e., the neutral state of $N^{\ast}$. It is natural to introduce an internal symmetry for the two fermions in a doublet, in particular an exchange symmetry between them, since there is no distinction between them a priori; in addition to the conventional chiral symmetry, the doublet can have an extra symmetry, either $\mathrm{U}(1)_{(1-2)\mathrm{L}}\times\mathrm{U}(1)_{(1-2)\mathrm{R}}$ (the relative phase of the doublet) or $\mathrm{SU}(2)_{\mathrm{L}}\times\mathrm{SU}(2)_{\mathrm{R}}$, dubbed an emergent chiral symmetry. The emergent chiral symmetries include the naive and mirror assignments as subgroups, thus providing us in-depth understanding of the chiral symmetry. We will construct the full Lagrangian including four-point interactions among the neutron $n$ and its chiral partner $n^{\ast}$ and present the phase structure in the mean field approximation. We find novel vector-type $J^{P}=1^{+}$ and $1^{-}$ condensates (${}^{3}P_{1}$ and ${}^{3}S_{1}$-${}^{3}D_{2}$ in nonrelativistic notation) exhibiting the dynamical breaking of the emergent chiral symmetries, the baryon number symmetry, and the rotational symmetry of the real space, resulting in emergent pions, superfluid phonons, and magnons, respectively as gapless Nambu-Goldstone (NG) bosons consisting of six quarks. We also find low-energy fermionic excitation modes exhibiting a spatial anisotropy of the propagation. We shall name those new states of superfluidity emergent chiral superfluids. The emergent chiral superfluids may exist in a more inner part of NSs than the phases of neutron ${}^{1}S_{0}$ and ${}^{3}P_{2}$ superfluids as illustrated in Fig. 1. It has been shown numerically that the neutron ${}^{3}P_{2}$ pairings can exist at baryon number densities up to $3n_{0}$ or a few more (see, e.g., Refs. [27, 28, 29]). Therefore, it should be expected that the baryon number densities relevant to the emergent chiral superfluids would be larger than those densities. [^2] In the present study, we consider the case that the critical density for the restoration of chiral symmetry should be higher than the density of ${}^{3}P_{2}$ superfluids.
@@
The paper is organized as follows: in Sec. II, we introduce the emergent chiral symmetry for the parity-doubled neutrons as an extension of the usual chiral symmetry. Based on the emergent chiral symmetry, we construct the Lagrangian with four-point interactions for the chiral-doublet baryons. In Sec. III, we adopt the Namb-Gor’kov formalism to the particle-particle (hole-hole) interaction in the parity-doubled baryons and apply the mean field approximation by introducing gap functions. We show that the vector-type condensate accompanies the breaking of the emergent chiral symmetry, baryon number symmetry, and rotational symmetry in the ground state. We also analyze the fermionic excitations in low-energy scale and show the existence of the Dirac points in the momentum space. The final section is devoted to our conclusion and outlook. In Appendix A, the naive and mirror assignments are briefly reviewed along with the emergent chiral symmetry. In Appendix B, the Lagrangians in the two assignments are given for the simplest emergent chiral symmetry. In Appendix C, the Fierz identities are summarized for convenience of readers. In Appendix D, the eigenfunctions of the Hamiltonian for the massless Dirac fermions are summarized.
@@
Figure 1: Schematic cross section inside a neutron star formed by the crust, the neutron ${}^{1}S_{0}$ superfluids near the surface, the neutron ${}^{3}P_{2}$ superfluids at intermediate region, and the emergent chiral superfluids in the deep interior. Phases of the emergent chiral superfluids are denoted by ${\cal S}_{a}1^{-}$ (${}^{3}P_{1}$) ($a=0,1,3$) and ${\cal A}1^{+}$ (${}^{3}S_{1}$-${}^{3}D_{1}$). Here, ${\cal S}_{a}$ and ${\cal A}$ indicate the symmetric and antisymmetric representations in the emergent chiral symmetry, and $J^{P}$ represents total angular momentum $J$ and parity $P$. In the parentheses, the nonrelativistic quantum numbers, corresponding to $J^{P}=1^{-}$ and $1^{+}$, are denoted by ${}^{2S+1}L_{J}$ with spin $S$ and angular momentum $L$.
@@ II.1 Naive and mirror assignments
To describe consistently a system for baryons both in the broken and restored phases of chiral symmetry, we introduce parity-doubled baryons which are composed of a pair of a nucleon $N$ with spin-parity $J^{P}=1/2^{+}$ and an excited nucleon $N^{\ast}(1535)$ with $J^{P}=1/2^{-}$ [77]. The parity-doublet model has been applied to the hadron spectroscopy related to the chiral dynamics, see, e.g., Ref. [13] for a review. In this framework, the nucleon $N$ and the excited nucleon $N^{\ast}(1535)$ are expressed as a linear combination of two baryons denoted by $\Psi_{1}$ and $\Psi_{2}$ in the flavor eigenstates. Assigning the chiralities to $\Psi_{1}$ and $\Psi_{2}$ ($N$ and $N^{\ast}$), two different types of chiral symmetry have been considered: a naive assignment and a mirror assignment [77]. In the naive assignment, the left (right) component of $\Psi_{1}$ and the left (right) component of $\Psi_{2}$ rotate simultaneously. In the mirror assignment, in contrast, the left (right) component of $\Psi_{1}$ and the right (left) component of $\Psi_{2}$ rotate simultaneously. Notice the different combinations of the left and right components for $\Psi_{1}$ and $\Psi_{2}$, respectively, in the mirror assignment.  [^3] A parity doublet has also been referred to as a chiral doublet in the literature, e.g. [78, 79, 80, 81].
To model a neutron superfluid in the context of parity doublet framework, we shall consider the interaction between two parity-doubled baryons, i.e., $N$ and $N^{\ast}$. Because our main interest is primarily put on the neutron stars, we consider only the neutral components $n$ and $n^{\ast}$ by neglecting the charged ones. Here $n^{\ast}$ indicates $n^{\ast}(1535)$, i.e., the neutral component of $N^{\ast}(1535)$, for shorthand notation. Thus, $n$ and $n^{\ast}$ are expressed as a linear combination of $\psi_{1}$ and $\psi_{2}$: $n=\cos\theta\psi_{1}+\gamma_{5}\sin\theta\psi_{2}$ and $n^{\ast}=-\gamma_{5}\sin\theta\psi_{1}+\cos\theta\psi_{2}$ with an appropriate mixing angle $\theta$ [12, 13]. Here we consider that $\psi_{1}$ and $\psi_{2}$ have only neutral components in $\Psi_{1}$ and $\Psi_{2}$, respectively. Characterizing our superfluid, we have $nn$, $nn^{\ast}$ and $n^{\ast}n^{\ast}$ interactions to get the particle-particle (hole-hole) pairings. We will consider a full set of four-point interactions (zero-range force) for $n$ and $n^{\ast}$. This setup is much different from the previous treatments of the parity-doublet model [6, 7, 8, 9, 10], where the $\sigma$ and $\omega$ fields are introduced to couple to the parity-doubled baryons like in the relativistic mean-field model of nuclear matter [82].